Clock and Cane Company has 8.00 percent, semiannual coupon bonds with face value of $1000 and twelve years left to maturity, trading in the market at a price of $900. Using this information, answer the following questions (a) – (c):

  1. What is the YTM of the bonds?
  1. What is the current yield on the bond?
  1. Suppose the last coupon payment was 2 months ago, what is the dirty/invoice price of the bond?
Answer #1

a) Using financial calculator to calculate ytm

Inputs: N= 12 × 2 = 24

Fv= 1,000

Pv= 900

Pmt= 8%/2 = 4% × 1,000 = 40

I/y = compute

We get, ytm of the bond as 4.70% × 2 = 9.40%

b) Current yield = Annual coupon payment/ current bond price

= 80 / 900

= 8.89%

C) Dirty price = clean price + accrued interest

Clean price= present value of future cash flow

= 900

Accrued interest= (Annual interest/ number of payment in year) × Time held after last coupon/time between two coupon

= 80 / 2 × 60/360

= 40 × 1/6

= 40/6

= $6.67

Dirty price = clean price + accrued interest

= 900 + 6.67

= $906.67